OpenStudy (anonymous):

A card is picked from a standard deck of cards. Determine the odds against and the odds in favor of selecting A card greater than 5 (ace is low).

OpenStudy (anonymous):

p(greater than 5) = 36/52 P(not greater than 5) = 16/52 so odd against : odd in favour = 36/52 : 16/52 = 36:16 = 9:4